OpenStudy (anonymous):

find length of polar curve. r=6/(1+cos(theta), 0<= theta<=(pi/2)

OpenStudy (anonymous):

\[r=6/1+cos \theta , 0\le \theta \le (\pi/2)\]

OpenStudy (anonymous):

\[\int\limits_{ }^{ }\sqrt{r ^{2} - (r \prime)^{2}} d \theta \]

OpenStudy (anonymous):

that second term is { dr/d(theta) }^2